´ e à é à éé à à é ê é à é é ê é é è à é é é à é à é é é è é è é éé é é éé ê éé é è é éé éé ê ô é à é é é é é é é é à é à é é à é à é è à é é é é è é à î è é è é

1

é ê à é é è à é é à é è è è è é éé é of_{u}niverse.jpg = notammentpourlecopyrightdel′image.RemerciementsàlaNASA.[width = 6cm]End_{o}f_{u}niverse.eps

é é é é é du_{s}oleil.jpg =

[width = 14cm]V ie_{d}u_{s}oleil.eps é ê è é ê è œ

Voir l’article Quelle est la forme de l’univers, Science et Vie junior, avril